OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be only 1 interaction. If there are 3 children, there can be 4 interactions. If there are 4 children, there can be 9 interactions. Which recursive equation represents the pattern?

OpenStudy (lgbasallote):

hint: 1 = 1^2 4 = 2^2 9 = 3^2 do you see the pattern?

OpenStudy (anonymous):

i dont understand the part on Which recursive equation represents the pattern?

OpenStudy (anonymous):

please help